# Hybrid Video Anomaly Detection for Anomalous Scenarios in Autonomous Driving

### v1

@@ Abstract
In autonomous driving, the most challenging scenarios are the ones that can only be detected within their temporal context. Most video anomaly detection approaches focus either on surveillance or traffic accidents, which are only a subfield of autonomous driving. In this work, we present HF2-VADAD, a variation of the HF2-VAD surveillance video anomaly detection method for autonomous driving. We learn a representation of normality from a vehicle’s ego perspective and evaluate pixel-wise anomaly detections in rare and critical scenarios.
@@
Keywords: Autonomous Driving, Anomaly Detection, Behavioral Anomalies
@@ 1 Introduction
In autonomous driving, the detection of anomalies is crucial for ensuring safety and reliability. Video anomaly detection (VAD) focuses on identifying events in video data that deviate from an expected normality. In autonomous driving, the challenges of detection are complicated by factors such as camera movements, ever-changing backgrounds, and rapid changes in vehicle speed.
Many different types of anomalies exist [1, 2, 3], with many approaches trying to detect them [4]. We can distinguish between five key techniques: Reconstruction, prediction, generative modeling, feature extraction, and confidence evaluation [5]. Reconstruction-based methods focus on learning a representation of normality and labeling deviations from this norm [6, 7]. Predictive approaches detect anomalies when future predictions deviate significantly from observations [8, 9]. However, in autonomous driving, most of the anomaly detection methods that are based on raw sensory data focus on atypical objects rather than atypical behaviors. Temporal anomalies are typically detected in object-based trajectory data, assuming perfect perception [10, 11, 12, 13]. The most complex temporal anomalies are anomalous scenarios, which pose a high risk of collision [1].
Figure 1: Exemplary anomaly detection with HF2-VADAD. In the left frame, the vehicle in front performs a rare sudden braking maneuver in simulation, as highlighted in yellow in the top-left graph. The graph shows the anomaly scores per frame, peaking while the vehicle performs the maneuver. The right frame shows pixel-wise anomaly scores for all instances of the class vehicle.
Until recently, mostly accident-focused benchmarks were available to evaluate temporal anomalies in the context of autonomous driving [14]. Given a rise in datasets with novel dynamic anomalies [15, 16, 17], we are interested in the applicability of video anomaly detection methods for autonomous driving. Among recent state-of-the-art methods, we choose HF2-VAD as the representative model [18] and adapt it to the field of autonomous driving. For this purpose, we train a HF2-VADAD model on simulated data representing normality and evaluate it on the challenging AnoVox dataset [17] with anomalous scenarios in the form of sudden braking maneuvers of the lead vehicle. We introduce a pixel-wise evaluation of HF2-VADAD to also localize anomalies.
@@ 2 Related Work
There are three typical camera perspectives for anomaly detection methods in the traffic domain. Birds-Eye-View (BEV) and surveillance perspectives come from external data providers, while an ego or dashcam view can be provided by a vehicle itself [14]. Numerous works have focused on specific scenarios, including the detection of atypical activity in urban surveillance, the analysis of irregularities in pedestrian dynamics, and the monitoring of abnormal behavior patterns in crowds [19, 20, 21, 22, 23, 24, 25, 26, 27] as well as street scenes recorded by city cameras that evaluate vehicle behavior from an outside perspective [28, 29].
Especially the VAD subfields traffic anomaly detection [30, 31] and road accident detection [32] have gained much traction. However, only a few methods in the literature specifically analyze behavioral anomalies from the perspective of a vehicle’s onboard cameras. For our comparison in Table 1, we have also excluded frame-wise [33, 34, 35, 36] or supervised [37, 38, 39, 40] methods as well as such that focus on domain-level anomalies [41, 42, 43, 44].
@@
In general video anomaly detection, hybrid methods that combine both reconstruction and prediction methods have emerged as the current state-of-the-art. These methods aim to utilize the strengths of both approaches by leveraging temporal dynamics and reconstruction capabilities to improve anomaly detection [45, 46, 18, 47, 48, 49]. Among these, HF2-VAD by Liu et al. [18], who built upon the work of Park et al. [50], is commonly used in the literature as a benchmark for the state of the art in video anomaly detection [51] and often serves as the foundation for further advancements [52]. It detects unusual events that deviate from normal patterns for surveillance tasks.
@@
The idea of predicting future frames for the purpose of anomaly detection became popular around 2019, where it was demonstrated both for surveillance tasks [8] and autonomous driving [53, 54]. Liu et al. combined future frame predictions with an optical flow [8], and Bolte et al. focused on semantically distinguishing relevant classes from the rest of an image [53]. Yao et al., on the other hand, focused on predicting only the bounding boxes of relevant classes in order to compensate for the motion of the ego vehicle, which makes the prediction of whole frames more complex [54]. Later, Fang et al. built upon their work but removed the need for an odometry estimation for the ego vehicle. Recently, Yao et al. extended their own work and also provided an updated version of Liu’s approach by adding masks for relevant objects, similar to how Bolte approached the topic [55]. Finally, the new DeepAccident dataset by Wang et al. also proposed an accident detection method based on future trajectories and their distances to each other [15]. As shown in Table 1, most of the methods require many additional sub-methods to function, making them rather complex and only recently began to work for high-resolution input images.
Table 1: Behavioral, pixel-wise anomaly detection methods for vehicle camera data in motion.
@@
AnoPred [8, 54, 55] 2019 Learned Normality
CC-Detector [53] 2019 Learned Normality
SSC-TAD [56] 2022 Learned Normality
FOL [55, 54] 2023 Learned Normality
AnoPredMask [8, 55] 2023 Learned Normality
AccPred [15] 2023 Metric-based
HF2-VADAD [18] 2024 Learned Normality
@@ 3 Methodology
As shown in Section 2 and Table 1, existing video anomaly detection methods for autonomous driving are either overly complex or incapable of dealing with high-resolution input data. Based on HF2-VAD [18], our adaption HF2-VADAD is capable of utilizing high-resolution video streams, learns a representation of normality in the domain of autonomous driving and only requires two sub-methods. Thus, it is the least complex method for video anomaly detection in the field of autonomous driving for high-resolution input.
Figure 2: HF2-VADAD: Adaptation and extension of HF2-VAD for autonomous driving. Optical flows ${y}_{1:t}$ and bounding boxes ${x}_{1:t}$ for relevant objects are generated for each input image. ML-MemAE-SC reconstructs the optical flows $\hat{y}_{1:t}$ with memory-modules $M$ to better reconstruct only normal patterns. The CVAE predicts a future frame $\hat{x}_{t+1}$. On this basis, image-wise and localized pixel-wise anomaly scores are generated. Adapted from [18].
HF2-VAD. Here, we provide a brief overview of the hybrid framework HF2-VAD [18] we build upon and refer to the original publication for further details. The framework combines Multi-Level Memory modules in an Autoencoder with Skip Connections (ML-MemAE-SC) for optical flow reconstructions and a Conditional Variational Autoencoder (CVAE) for future frame predictions. The framework is trained on normal data and uses discrepancies in the reconstructions and predictions to identify anomalies. Experiments were conducted with well-established surveillance datasets.
@@
The CVAE plays a central role in predicting future images by using conditioned historical data to simulate probable future scenarios. It determines the latent variable $z_{t+1}$ by considering both the current image $x_{t}$ and the optical flow $y_{t,t+1}$ between successive images. It works on the premise that a latent variable $z$, representing the probabilistic state of the following image, can be derived from the immediate past to subsequently influence the generation of $\hat{x}_{t+1}$. It uses the encoded optical flow and the current image to estimate the posterior distribution $p(z_{t+1}|x_{t},y_{t,t+1})$. A sample from this distribution, which represents the probabilistic state of the next image, is then drawn for image generation. During training, this latent variable $z$ is combined with the conditions derived from the encoder $E$ and processed by the decoder $D$ to create the future image $\hat{x}_{t+1}$.
@@
HF2-VADAD. First, we adapt HF2-VAD to autonomous driving by generating dense, pixel-wise anomaly scores for the whole frame rather than focusing on the content of unlocalized, detected bounding boxes alone, as can be seen in Figure 1. We learn flow reconstructions and flow-guided frame prediction from the perspective of an ego-vehicle rather than from the perspective of a static surveillance camera. Our training dataset was collected with the CARLA simulation engine [57]. Rather than collecting data with the default autopilot, we employ Roach [58] as the driving agent.
@@
Flows are generated by FlowNet2.0  [59] to ensure that both static and motion-based features contribute to the training of the model. Bounding boxes are detected with an off-the-shelf object detection model [60] trained on CARLA data.
We first train the ML-MemAE-SC with the extracted flows from the training dataset, then train the CVAE model with the reconstructed flows from the ML-MemAE-SC and the corresponding frames, and finally fine-tune the whole framework. During fine-tuning, the weights of the two models are adjusted together to improve their coordination and the overall accuracy of anomaly detection.
@@
Figure 3: Sequential Anomaly Detection: The first image shows the normal driving mode of the lead vehicle, while the last two images show an anomaly with the pixel-wise ground truth overlayed in red. HF2-VADAD successfully detects the unknown maneuver.
In addition to the existing frame-wise anomaly detection, we have integrated a pixel-wise anomaly detection. For this, the anomaly score for each pixel, denoted by $S_{\text{pixel}}$, is derived from a weighted average of robustly scaled mean squared errors (MSE) for optical flow and image prediction:
$S_{\text{pixel}}=w_{\text{r,p}}\cdot m_{\text{r}}+w_{\text{p,p}}\cdot m_{\text{p}}$ (2)
The variables $m_{\text{r}}$ and $m_{\text{p}}$ denote the mean values of the robustly scaled MSEs for the optical flow and image prediction, respectively. The scores are calculated using the weights $w_{\text{r,p}}$ for optical flow and $w_{\text{p,p}}$ for image prediction. Robust scaling is used to standardize the MSEs by reducing the influence of outliers, making the method more reliable for detecting anomalies under different conditions. The averaging process calculates the mean of these scaled values to summarize the error across all pixels. The resulting pixel-wise scores are then used to localize anomalies within the image. A score of 0 is set for all pixels outside the bounding boxes, as shown in Figure 1. In contrast, the frame-wise approach of the original HF2-VAD uses simple summation and normalization based on the training statistics. Specifically, the sum of MSE losses across all dimensions is calculated for both optical flow and image predictions, and these summed values are then normalized using the mean and standard deviation from the training data. In the presence of detections, they add up all $32\times 32$ anomaly patches and choose the largest value as the frame-wise anomaly score. We also report these frame-wise results as AUROC in Section 4.
In contrast, our approach calculates a pixel-wise anomaly score by normalizing and aggregating mean-squared errors of optical flow and frame predictions, then scaling these scores to bounding boxes within each frame. This pixel-wise approach allows for a more granular analysis, focusing on specific areas within frames rather than a single aggregated frame-level score.

### v2

# Karlsruhe, Germany Hybrid Video Anomaly Detection for AD

@@ Abstract
In autonomous driving, the most challenging scenarios can only be detected within their temporal context. Most video anomaly detection approaches focus either on surveillance or traffic accidents, which are only a subfield of autonomous driving. We present HF2-VADAD, a variation of the HF2-VAD surveillance video anomaly detection method for autonomous driving. We learn a representation of normality from a vehicle’s ego perspective and evaluate pixel-wise anomaly detections in rare and critical scenarios.
@@ 1 Introduction
In autonomous driving, the detection of anomalies is crucial for ensuring safety and reliability. Video anomaly detection (VAD) focuses on events in video data that deviate from an expected normality. In autonomous driving, the challenges are complicated by factors such as camera movements, ever-changing backgrounds, and rapid changes in vehicle speed.
Many different types of anomalies exist [Breitenstein et al.(2020)Breitenstein, Termöhlen, Lipinski, and Fingscheidt, Heidecker et al.(2021)Heidecker, Breitenstein, Rösch, Löhdefink, Bieshaar, Stiller, Fingscheidt, and Sick, Bogdoll et al.(2021)Bogdoll, Breitenstein, Heidecker, Bieshaar, Sick, Fingscheidt, and Zollner], with many approaches trying to detect them [Bogdoll et al.(2022)Bogdoll, Nitsche, and Zöllner]. Reconstruction-based methods focus on learning a representation of normality and labeling deviations from this norm [Hasan et al.(2016)Hasan, Choi, Neumann, Roy-Chowdhury, and Davis, Luo et al.(2017)Luo, Liu, and Gao]. Predictive approaches detect anomalies when future predictions deviate significantly from observations [Liu et al.(2018)Liu, Luo, Lian, and Gao, Yu et al.(2020)Yu, Wang, Cai, Zhu, Xu, Yin, and Kloft]. However, in autonomous driving, most of the anomaly detection methods that are based on raw sensory data focus on atypical objects rather than atypical behaviors. Temporal anomalies are often detected in object-based trajectory data, assuming perfect perception [Wiederer et al.(2021)Wiederer, Bouazizi, Troina, Kressel, and Belagiannis, Jiao et al.(2023)Jiao, Bai, Liu, Sato, Yuan, Chen, and Zhu, Breuer et al.(2021)Breuer, Le Xuan, Termöhlen, Homoceanu, and Fingscheidt, Hsu et al.(2023)Hsu, Leung, Chen, Fisac, and Pavone]. The most complex temporal anomalies are anomalous scenarios, which pose a high risk of collision [Breitenstein et al.(2020)Breitenstein, Termöhlen, Lipinski, and Fingscheidt].
Figure 1: Exemplary anomaly detection with HF2-VADAD. The vehicle in front performs a sudden braking maneuver, as highlighted in yellow in the top-left graph. The graph shows the anomaly scores per frame, peaking while the vehicle performs the maneuver. The right frame shows pixel-wise anomaly scores for all instances of the class vehicle, computing anomaly scores only within class-specific bounding boxes predicted by an object detector.
Until recently, mostly accident-focused benchmarks were available to evaluate temporal anomalies in the context of autonomous driving [Fang et al.(2024b)Fang, Qiao, Xue, and Li]. Given a rise in datasets with novel dynamic anomalies [Wang et al.(2024)Wang, Kim, Wenxuan, Xie, Ge, Chen, Li, and Luo, Gong et al.(2024)Gong, Zhang, Xia, Zhang, and Ji, Bogdoll et al.(2024)Bogdoll, Hamdard, Rößler, Geisler, Bayram, Wang, Imhof, de Campos, Tabarov, Yang, Gottschalk, and Zöllner], we are interested in the applicability of video anomaly detection methods for autonomous driving. Among recent SotA methods, we choose HF2-VAD as the representative model [Liu et al.(2021)Liu, Nie, Long, Zhang, and Li] and adapt it to the field of autonomous driving. For this purpose, we train a HF2-VADAD model on simulated data representing normality and evaluate it on the AnoVox benchmark [Bogdoll et al.(2024)Bogdoll, Hamdard, Rößler, Geisler, Bayram, Wang, Imhof, de Campos, Tabarov, Yang, Gottschalk, and Zöllner] with anomalous scenarios in the form of sudden braking maneuvers of a lead vehicle. We also introduce a pixel-wise evaluation to localize anomalies.
@@ 2 Related Work
There are three typical camera perspectives for anomaly detection methods in the traffic domain. Birds-Eye-View (BEV) and surveillance perspectives come from external data providers, while an ego or dashcam view can be provided by a vehicle itself [Fang et al.(2024b)Fang, Qiao, Xue, and Li]. Numerous works have focused on specific scenarios, including the detection of atypical activity in urban surveillance, the analysis of irregularities in pedestrian dynamics, and the monitoring of abnormal behavior patterns in crowds [Chandola et al.(2009)Chandola, Banerjee, and Kumar, Popoola and Kejun Wang(2012), Nayak et al.(2021)Nayak, Pati, and Das, Ramachandra et al.(2020)Ramachandra, Jones, and Vatsavai, Tran et al.(2022)Tran, Vu, Vo, Nguyen, and Nguyen, Santhosh et al.(2020)Santhosh, Dogra, and Roy, Aldayri and Albattah(2022), Patil and Biswas(2016), Chandrakala et al.(2023)Chandrakala, Deepak, and Revathy] as well as street scenes recorded by city cameras that evaluate vehicle behavior from an outside perspective [Kumaran et al.(2021)Kumaran, Dogra, and Roy, Djenouri et al.(2019)Djenouri, Belhadi, Lin, Djenouri, and Cano].
@@
In general video anomaly detection, hybrid methods that combine both reconstruction and prediction methods have emerged as the current state-of-the-art. These methods aim to utilize the strengths of both approaches by leveraging temporal dynamics and reconstruction capabilities to improve anomaly detection [Ye et al.(2019)Ye, Peng, Gan, Wu, and Qiao, Tang et al.(2020)Tang, Zhao, Zhang, Gong, Li, and Yang, Liu et al.(2021)Liu, Nie, Long, Zhang, and Li, Liu et al.(2022)Liu, Zhang, Niu, and Wang, Li et al.(2023)Li, Yang, Shen, Zhang, and Long, Huang et al.(2023)Huang, Zhao, Gao, Chen, Wang, and Hussain]. Among these, HF2-VAD by Liu et al. [Liu et al.(2021)Liu, Nie, Long, Zhang, and Li], who built upon the work of Park et al. [Park et al.(2020)Park, Noh, and Ham], is commonly used in the literature as a benchmark for the state of the art in video anomaly detection [Ristea et al.(2024)Ristea, Croitoru, Ionescu, Popescu, Khan, and Shah] and often serves as the foundation for further advancements [Wan et al.(2023)Wan, Zhang, and Jin]. It detects unusual events that deviate from normal patterns for surveillance tasks.
@@
The idea of predicting future frames for the purpose of anomaly detection became popular around 2019, where it was demonstrated both for surveillance tasks [Liu et al.(2018)Liu, Luo, Lian, and Gao] and autonomous driving [Bolte et al.(2019)Bolte, Bar, Lipinski, and Fingscheidt, Yao et al.(2019)Yao, Xu, Wang, Crandall, and Atkins]. Liu et al. combined future frame predictions with an optical flow [Liu et al.(2018)Liu, Luo, Lian, and Gao] in AnoPred, and CC-Detector by Bolte et al. focused on semantically distinguishing relevant classes from the rest of an image [Bolte et al.(2019)Bolte, Bar, Lipinski, and Fingscheidt]. Fang et al. proposed self-supervised consistency learning for traffic accident detection (SSC-TAD[Fang et al.(2022)Fang, Qiao, Bai, Yu, and Xue], where the key idea is to detect deviations of appearance, motion, and context relations. For this, they also utilize an explicit graph representation of the scenes. Future object localization (FOL) by Yao et al., on the other hand, focused on predicting only the bounding boxes of relevant classes in order to compensate for the motion of the ego vehicle, which makes the prediction of whole frames more complex [Yao et al.(2019)Yao, Xu, Wang, Crandall, and Atkins]. Later, Fang et al. built upon their work but removed the need for an odometry estimation for the ego vehicle. Recently, Yao et al. extended their own work and also provided updated AnoPredMask version of Liu’s approach by adding masks for relevant objects, similar to Bolte’s approach [Yao et al.(2023)Yao, Wang, Xu, Pu, Wang, Atkins, and Crandall]. Finally, the new DeepAccident dataset by Wang et al. proposed AccPred, an accident detection method based on future trajectories and their distances to each other [Wang et al.(2024)Wang, Kim, Wenxuan, Xie, Ge, Chen, Li, and Luo]. As shown in Table 1, most of the methods require many additional sub-methods to function, making them rather complex and only recently began to work for high-resolution input images.
@@
AnoPred [Liu et al.(2018)Liu, Luo, Lian, and Gao, Yao et al.(2019)Yao, Xu, Wang, Crandall, and Atkins, Yao et al.(2023)Yao, Wang, Xu, Pu, Wang, Atkins, and Crandall] 2019 Learned Normality
CC-Detector [Bolte et al.(2019)Bolte, Bar, Lipinski, and Fingscheidt] 2019 Learned Normality
SSC-TAD [Fang et al.(2022)Fang, Qiao, Bai, Yu, and Xue] 2022 Learned Normality
FOL [Yao et al.(2023)Yao, Wang, Xu, Pu, Wang, Atkins, and Crandall, Yao et al.(2019)Yao, Xu, Wang, Crandall, and Atkins] 2023 Learned Normality
AnoPredMask [Liu et al.(2018)Liu, Luo, Lian, and Gao, Yao et al.(2023)Yao, Wang, Xu, Pu, Wang, Atkins, and Crandall] 2023 Learned Normality
AccPred [Wang et al.(2024)Wang, Kim, Wenxuan, Xie, Ge, Chen, Li, and Luo] 2023 Metric-based
HF2-VADAD [Liu et al.(2021)Liu, Nie, Long, Zhang, and Li] 2024 Learned Normality
Table 1: Behavioral, pixel-wise anomaly detection methods for vehicle camera data.
@@ 3 Methodology
As shown in Section 2 and Table 1, existing video anomaly detection methods for autonomous driving are either overly complex or incapable of dealing with high-resolution input data. Based on HF2-VAD [Liu et al.(2021)Liu, Nie, Long, Zhang, and Li], our adaption HF2-VADAD is capable of utilizing high-resolution video streams, learns a representation of normality in the domain of autonomous driving and only requires two sub-methods. Thus, it is the least complex method for video anomaly detection in the field of autonomous driving for high-resolution input.
Figure 2: HF2-VADAD: Adaptation and extension of HF2-VAD for autonomous driving. Optical flows ${y}_{1:t}$ and bounding boxes ${x}_{1:t}$ for relevant objects are generated for each input image. ML-MemAE-SC reconstructs the optical flows $\hat{y}_{1:t}$ with memory-modules $M$ to better reconstruct only normal patterns. The CVAE predicts a future frame $\hat{x}_{t+1}$. On this basis, image-wise and localized pixel-wise anomaly scores are generated. Adapted from [Liu et al.(2021)Liu, Nie, Long, Zhang, and Li].
HF2-VAD. Here, we provide a brief overview of the hybrid framework HF2-VAD [Liu et al.(2021)Liu, Nie, Long, Zhang, and Li] we build upon and refer to the original publication for further details. The framework combines Multi-Level Memory modules in an Autoencoder with Skip Connections (ML-MemAE-SC) for optical flow reconstructions and a Conditional Variational Autoencoder (CVAE) for future frame predictions. The framework is trained on normal data and uses discrepancies in the reconstructions and predictions to identify anomalies. Experiments were conducted with well-established surveillance datasets.
@@
The CVAE plays a central role in predicting future images by using conditioned historical data to simulate probable future scenarios. It determines the latent variable $z_{t+1}$ by considering both the current image $x_{t}$ and the optical flow $y_{t,t+1}$ between successive images. It works on the premise that a latent variable $z$, representing the probabilistic state of the following image, can be derived from the immediate past to subsequently influence the generation of $\hat{x}_{t+1}$. It uses the encoded optical flow and the current image to estimate the posterior distribution $p(z_{t+1}\mid x_{t},y_{t,t+1})$. A sample from this distribution, which represents the probabilistic state of the next image, is then drawn for image generation. During training, this latent variable $z$ is combined with the conditions derived from the encoder $E$ and processed by the decoder $D$ to create the future image $\hat{x}_{t+1}$.
@@
HF2-VADAD. As shown in Figure 2, we first adapt HF2-VAD to autonomous driving by generating dense, pixel-wise anomaly scores for the whole frame rather than focusing on the content of unlocalized, detected bounding boxes alone, as can be seen in Figure 1. We learn flow reconstructions and flow-guided frame prediction from the perspective of an ego-vehicle rather than from the perspective of a static surveillance camera. Our training dataset was collected with the CARLA simulation engine [Dosovitskiy et al.(2017)Dosovitskiy, Ros, Codevilla, Lopez, and Koltun]. Rather than collecting data with the default autopilot, we employ Roach [Zhang et al.(2021)Zhang, Liniger, Dai, Yu, and Van Gool] as the driving agent.
@@
Flows are generated by FlowNet2.0 [Ilg et al.(2017)Ilg, Mayer, Saikia, Keuper, Dosovitskiy, and Brox] to ensure that both static and motion-based features contribute to the training of the model. Bounding boxes are detected with an off-the-shelf object detection model [MinyiLin(2022)] trained on CARLA data. We first train the ML-MemAE-SC with the extracted flows from the training dataset, then train the CVAE model with the reconstructed flows from the ML-MemAE-SC and the corresponding frames, and finally fine-tune the whole framework. During fine-tuning, the weights of the two models are adjusted together to improve their coordination and the overall accuracy of anomaly detection.
@@
Figure 3: Sequential Anomaly Detection: The first image shows the normal driving mode of the lead vehicle, while the last two images show an anomaly with the pixel-wise ground truth overlayed in red. Other regularly driving traffic participants are also present. As visible in the anomaly maps, HF2-VADAD successfully detects the unknown maneuver.
In addition to frame-wise anomaly detection, we have integrated a pixel-wise anomaly detection. The pixel-wise anomaly score, denoted as $S_{\text{pixel}}$, is computed by combining the MSE values from optical flow and image prediction, both of which are normalized to enhance the reliability of the anomaly detection process:
$S_{\text{pixel}}=w_{\text{r,p}}\cdot m_{\text{r}}+w_{\text{p,p}}\cdot m_{\text{p}}$
Here, $m_{\text{r}}$ and $m_{\text{p}}$ represent the aggregated robustly scaled MSE values for optical flow and image prediction, respectively. To enhance the robustness and comparability of these MSE values across different datasets, a robust scaling technique is applied. This scaling is important as it mitigates the influence of outliers and non-standard data distributions, thereby improving the reliability of the anomaly detection process. The robust scaling of each MSE value x is defined as follows:
$x_{\text{scaled}}=\frac{x-\tilde{x}}{\text{IQR}(X)}$
Where $x$ is the original MSE value, $\tilde{x}$ is the median of the MSE values across the dataset, and $\text{IQR}(X)$ is the interquartile range, defined as the difference between the 75th and 25th percentiles ($Q_{75}(X)$ and $Q_{25}(X)$). After robust scaling, the mean value is calculated for each pixel across the RGB channels:
$m_{\text{aggregated}}=\frac{1}{N}\sum_{i=1}^{N}x_{\text{scaled},i}$
Where $N$ is the number of RGB channels. The final pixel-wise anomaly score $S_{\text{pixel}}$ is computed by combining the aggregated MSE values from optical flow and image prediction using the weights $w_{\text{r,p}}$ and $w_{\text{p,p}}$.
The resulting pixel-wise scores are then used to localize anomalies within the image. A score of 0 is set for all pixels outside the bounding boxes, as shown in Figure 1. In contrast, the frame-wise approach of the original HF2-VAD uses simple summation and normalization based on the training statistics. Specifically, the sum of MSE losses across all dimensions is calculated for both optical flow and image predictions, and these summed values are then normalized using the mean and standard deviation from the training data. In the presence of detections, the authors added up all $32\times 32$ anomaly patches and chose the largest value as the frame-wise anomaly score. We report these frame-wise results as AUROC in Section 4.
In contrast, our approach calculates a pixel-wise anomaly score by normalizing and aggregating mean-squared errors of optical flow and frame predictions, then scaling these scores to bounding boxes within each frame. Pixel-wise scores allow for a localization of anomalies for a more granular analysis, compared to a single aggregated score per frame.